Nitrogen gas reacts with hydrogen gas to produce ammonia (NH3) gas. How many liters of NH3 can be produced at 0.93 atm and 24°C from a 16.0-g sample of nitrogen gas and an excess of hydrogen gas?

N2(g) + 3H2(g) => 2NH3(g)


1
Expert's answer
2021-11-15T10:58:56-0500

PV=nRT

Molar mass of nitrogen= 14.0067

16/14.0067=1.14

0.93×V=1.14×8.314×297

0.93V=2814.95

C=3026.83

=3.027L